OpenStudy (mimiii3):

What is the standard form for 2.83 x 10-4?

rebeccaxhawaii (rebeccaxhawaii):



Directrix (directrix):

2.83 x 10^(-4)= 0 0 0 0 0 0 0 0 . 2 8 3 * 10 *(-4)= 10^(-4) --> Move the decimal point to the left 4 times. 0 0 0 0 0 . 0 0 0 2. 8 3 or . 0 0 0 2. 8 3 which is .000283

Directrix (directrix):

@mimiii3